4,065.014 births in a year. 2,081,287 were boys, 1,983,727 were girls, 2 random women selected what is probability at least one child is a boy?
Answer #1

ANSWER:_

Probabilty of one chld boy =P(B1G2 +G1B2+B1B2) =0.512*0.488+0.488*0.512+0.512*0.512=0.761856

hence probabilty of other child boy given first one is boy =0.512*0.512/0.761856 =0.344086A